### v3

@@ 8 Conclusion
where $\mathcal{N}(i,-1)\equiv\varnothing$. As $r$ increases, we move away from the boundary and towards the interior. If $i\in J(r,C_{j})$, then $\mathcal{N}(i,r)\subseteq C_{j}$, so for such $i$, by Assumptions 3 and 4,
@@
$\leq 2c\,\max\{r,1\}^{-\gamma}.$
Using this calculation, for any $t\in\{0,1\}$,
@@
$\precsim\frac{1}{n}\sum_{j=1}^{k_{n}}\left(\lvert\mathcal{B}(C_{j})\rvert+\sum_{r=0}^{L_{n}}\sum_{i\in J(r,C_{j})}\max\{r,1\}^{-\gamma}\right)$
$\precsim\underbrace{\frac{k_{n}}{n}\xi_{n}\left(\frac{n}{k_{n}\xi_{n}}\right)^{\frac{d-1}{d}}}_{\text{boundary bias}}+\frac{k_{n}}{n}\sum_{r=0}^{L_{n}}\xi_{n}\max\{(L_{n}-r-1)^{d-1},1\}\max\{r,1\}^{-\gamma}$
$\precsim(k_{n}\xi_{n}/n)^{1/d}+\frac{k_{n}\xi_{n}}{n}L_{n}^{d-1}\sum_{r=1}^{\infty}r^{-\gamma}\precsim(k_{n}\xi_{n}/n)^{1/d}.$
@@
Fix any sequence $\{\xi_{n}\}_{n\in\mathbb{N}}$ such that $1\precsim\xi_{n}\prec n$. Let $\rho$ be the sup norm and $\mathcal{N}_{n}=\{\xi_{n}^{-1/d}x\colon x\in\mathbb{Z}^{d}\}\cap B(\bm{0},\mathcal{R}_{n})$ for $\mathcal{R}_{n}\in\mathbb{Z}$ and $\mathcal{R}_{n}\sim(n/\xi_{n})^{1/d}$. When $\xi_{n}\sim 1$, we have a cube $B(\bm{0},\mathcal{R}_{n})$ with radius $\sim n^{1/d}$ containing $\sim n$ units positioned on the integer lattice. When $\xi_{n}\succ 1$, we shrink all locations in the integer lattice and the radius of the cube towards the origin by a factor $\xi_{n}^{-1/d}$. This results in the same asymptotic order of number of units $\lvert\mathcal{N}_{n}\rvert\sim n$, but the volume of the cube is reduced to $\sim\mathcal{R}_{n}^{d}\sim n/\xi_{n}$, resulting in a density of $\xi_{n}$.
@@
$\lvert\mathcal{N}(i,r+\xi_{n}^{-1/d})\backslash\mathcal{N}(i,r)\rvert\in[\min\{C_{1}^{-1}\xi_{n}^{(d-1)/d}r^{d-1},n\},\max\{C_{1}\xi_{n}^{(d-1)/d}r^{d-1},1\}],$ (SA.6.17)
which implies
@@
$\leq 1+C_{1}d^{-1}\xi_{n}((r+\xi_{n}^{-1/d}[\xi_{n}^{1/d}])^{d}-r^{d}),$
where $[c]$ means round $c$ to the nearest integer. This verifies Assumption 2.
For the remainder of the proof, suppose the clusters are generated by $k$-medoids which by Theorem 1 satisfy Assumption 5.
@@
$Y_{i}(\bm{d})=d_{i}\frac{\sum_{j\in\mathcal{N}_{n}}d_{j}\bm{1}\{\rho(i,j)=\xi_{n}^{-1/d}\}}{\sum_{j\in\mathcal{N}_{n}}\bm{1}\{\rho(i,j)=\xi_{n}^{-1/d}\}},$
which is unit $i$’s treatment times the fraction of treated units in $i$’s “1-neighborhood” (so called because for $\xi_{n}=1$, these are the units at distance one from $i$). The sum in the denominator is always at least one by construction of $\mathcal{N}_{n}$, so Assumption 1 holds. Assumption 3 holds because potential outcomes only depend on “1-neighborhood” treatments.
Reduction to HT. Since $p_{0}=0$, $Y_{i}(\bm{0})=0$. Since $Q\in\{D,T,O\}$ and $p_{0}=0$, $T_{0i}^{+}Y_{i}=0$. Then
@@
Bias. Let $\beta=(p_{1}-p_{1}q)\bm{1}\{d_{1}=1\}+p_{1}(p_{1}-p_{1}q)\bm{1}\{d_{1}=\varnothing\}$, which is strictly positive under the assumptions of the theorem. We have
$\lvert{\bf E}[\tilde{\theta}^{+}]-\theta^{*}\rvert=\bigg{|}\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}\big{(}{\bf E}[Y_{i}\mid T_{1i}^{+}=1]-{\bf E}^{*}_{p_{1}}[Y_{i}(d_{1},\bm{D}_{-i})]\big{)}\bigg{|}\\ =\beta\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}\frac{\sum_{j\in\mathcal{N}_{n}\backslash C_{c(i)}}\bm{1}\{\rho(i,j)=\xi_{n}^{-1/d}\}}{\sum_{j\in\mathcal{N}_{n}}\bm{1}\{\rho(i,j)=\xi_{n}^{-1/d}\}}.$ (SA.6.18)
@@
~$\tilde{\mathcal{B}}(C)=\big{\{}i\in C\colon\lvert\mathcal{N}(i,\xi_{n}^{-1/d})\cap\mathbb{Z}^{d}\backslash C\rvert\geq 1\big{\}},$
the set of units whose “1-neighborhoods” intersect a point outside the cluster. Then
(SA.6.18)$\geq\beta\frac{1}{n}\sum_{j=1}^{k_{n}}\sum_{i\in\tilde{\mathcal{B}}(C_{j})}\frac{\sum_{\ell\in\mathcal{N}_{n}\backslash C_{j}}\bm{1}\{\rho(i,\ell)=\xi_{n}^{-1/d}\}}{\sum_{\ell\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell)=\xi_{n}^{-1/d}\}}.$ (SA.6.19)
By construction of $\mathcal{N}_{n}$ and the clusters,
$\min_{j=1,\ldots,k_{n}}\min_{i\in\tilde{\mathcal{B}}(C_{j})\backslash\tilde{\mathcal{B}}(\mathcal{N}_{n})}\frac{\sum_{\ell\in\mathcal{N}_{n}\backslash C_{j}}\bm{1}\{\rho(i,\ell)=\xi_{n}^{-1/d}\}}{\sum_{\ell\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell)=\xi_{n}^{-1/d}\}}\geq\min_{i\in\mathcal{N}_{n}}\left(\sum_{\ell\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell)=\xi_{n}^{-1/d}\}\right)^{-1}\succsim 1$
@@
Recalling the definitions in Assumption 5, let $m_{j}$ denote the “centroid” of a cluster $C_{j}$. Part (a) of the assumption implies that units in ~$\tilde{\mathcal{B}}(C_{j})$ are at least $\lfloor L_{n}\rfloor$ from $m_{j}$. For $r$ equal to multiples of $\xi_{n}^{1/d}$, define the “contour sets”
$\tilde{J}(r,C_{j})=\mathcal{N}(m_{j},r)\backslash\mathcal{N}(m_{j},r-\xi_{n}^{-1/d}).$
@@
$\sum_{\ell\in\mathcal{N}_{n}}\sum_{m\in\mathcal{N}_{n}}\frac{\bm{1}\{\rho(i,\ell)=\xi_{n}^{-1/d}\}}{\sum_{\ell^{\prime}\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell^{\prime})=\xi_{n}^{-1/d}\}}\frac{\bm{1}\{\rho(j,m)=\xi_{n}^{-1/d}\}}{\sum_{m^{\prime}\in\mathcal{N}_{n}}\bm{1}\{\rho(j,m^{\prime})=\xi_{n}^{-1/d}\}}\\ \times p_{1}^{-2}\text{Cov}\big{(}T_{1i}^{+}D_{i}D_{\ell},T_{1j}^{+}D_{j}D_{m}\big{)}.$
@@
$\text{Cov}(W_{c(i)}D_{i}D_{\ell},W_{c(j)}D_{j}D_{m})\\ =\underbrace{{\bf E}[\text{Cov}(D_{i}D_{\ell},D_{j}D_{m}\mid W_{c(i)})]}_{\geq 0}+\text{Cov}({\bf E}[D_{i}D_{\ell}\mid W_{c(i)}],{\bf E}[D_{j}D_{m}\mid W_{c(i)}]).$
@@
$\text{Var}\left(\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}Z_{i}^{+}\right)\geq\alpha\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in C_{c(i)}}\sum_{\ell\in C_{c(i)}\backslash\{i\}}\frac{\bm{1}\{\rho(i,\ell)=\xi_{n}^{-1/d}\}}{\sum_{\ell^{\prime}\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell^{\prime})=\xi_{n}^{-1/d}\}}\\ \times\sum_{m\in C_{c(i)}\backslash\{j\}}\frac{\bm{1}\{\rho(j,m)=\xi_{n}^{-1/d}\}}{\sum_{m^{\prime}\in\mathcal{N}_{n}}\bm{1}\{\rho(j,m^{\prime})=\xi_{n}^{-1/d}\}}\succsim\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\lvert C_{c(i)}\rvert.$
@@
By Assumption 1 and Lemma SA.5.3, there exists $c_{n}\precsim 1$ such that for all $i$,
$p_{ti}^{-1}T_{ti}\lvert Y_{i}(\bm{D})-{\bf E}[Y_{i}(\bm{D})\mid\mathcal{F}_{i}(r)]\rvert\leq 2c_{n}\min\{r^{-\gamma},1\}.$ (SA.6.23)
Hence, for any $q>0$,
${\bf E}[\lvert B_{i}\rvert^{q}]^{1/q}={\bf E}[\lvert Z_{i}-{\bf E}[Z_{i}\mid\mathcal{F}_{i}(r_{n})]\rvert^{q}]^{1/q}\leq 4c_{n}r_{n}^{-\gamma}.$ (SA.6.24)
By Lemma SA.5.4, $\max_{i}\lvert\Lambda_{i}\rvert\precsim n/k_{n}$, so using (SA.6.11),
$\lvert[P1]\rvert\leq\max_{i,j}{\bf E}[B_{i}^{2}]^{1/2}{\bf E}[B_{j}^{2}]^{1/2}\leq(4c_{n}r_{n}^{-\gamma})^{2}\precsim(k_{n}\xi_{n}/n)^{2\gamma/d},$
@@
Turning to $\lvert[P2]\rvert$, fix $i,j$ such that $j\not\in\Lambda_{i}$. Using (SA.6.24), there exists $c_{n}^{\prime}\precsim 1$ such that for all such $i,j$,
$\lvert{\bf E}[B_{i}B_{j}]\rvert\leq c_{n}^{\prime}r_{n}^{-\gamma}.$
@@
$B_{i}-\underbrace{{\bf E}[B_{i}\mid\mathcal{F}_{i}(s/2-\bar{R})]}_{X_{i}}=Z_{i}-{\bf E}[Z_{i}\mid\mathcal{F}_{i}(s/2-\bar{R})]$
since $\mathcal{N}(i,s/2-\bar{R})\supseteq\mathcal{N}(i,r_{n})$. Moreover, under Assumption 4, $X_{i}\perp\!\!\!\perp X_{j}$ since $\mathcal{N}(i,s/2-\bar{R})$ and $\mathcal{N}(j,s/2-\bar{R}))>2\bar{R}$ are separated by a distance of at least $2\bar{R}$, which is an upper bound on the “diameter” of any cluster. Applying (SA.6.24) with $s/2-\bar{R}$ in place of $r_{n}$, there exists $c_{n}^{\prime}\precsim 1$ such that for any $i,j$,
$\lvert{\bf E}[B_{i}B_{j}]\rvert\leq\lvert{\bf E}[X_{i}X_{j}]\rvert+\lvert{\bf E}[X_{i}(B_{j}-X_{j})]\rvert\\ +\lvert{\bf E}[(B_{i}-X_{i})X_{j}]\rvert+\lvert{\bf E}[(B_{i}-X_{i})(B_{j}-X_{j})]\rvert\leq c_{n}^{\prime}(s/2-\bar{R})^{-\gamma}.$
@@
$\lvert[P2]\rvert\leq\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\not\in\Lambda_{i}}\lvert{\bf E}[B_{i}B_{j}]\rvert\\ \leq k_{n}\frac{c_{n}^{\prime}}{n^{2}}\sum_{\ell=1}^{k_{n}}\sum_{i\in C_{\ell}}\sum_{s=\lfloor 2r_{n}\rfloor}^{\infty}\sum_{j\not\in\Lambda_{i}}\bm{1}\{\rho(i,j)\in[s-1,s)\}\\ \times\big{(}r_{n}^{-\gamma}\bm{1}\{s\leq 4\bar{R}\}+(s/2-\bar{R})^{-\gamma}\bm{1}\{s>4\bar{R}\}\big{)}.$
@@
$\lvert\sigma_{n}-\tilde{\sigma}_{n}\rvert\leq\text{Var}\left(\sqrt{k_{n}}\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}\mid\mathcal{F}_{i}(r_{n})])\right)^{1/2}\prec 1$ (SA.6.25)
@@
$\tilde{\sigma}_{n}^{-1}\sqrt{k_{n}}\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}\big{(}{\bf E}[Z_{i}\mid\mathcal{F}_{i}(r_{n})]-{\bf E}[Z_{i}]\big{)}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,1)$ (SA.6.26)
@@
so (SA.6.26) follows from his (3.8). This completes the proof of (6)
@@
$\sqrt{k_{n}}(\bar{\theta}-\theta^{*})\precsim\sqrt{k_{n}}(k_{n}\xi_{n}/n)^{\gamma/d}\prec\sqrt{k_{n}}(n/\xi_{n})^{-\frac{\gamma}{2\gamma+d}}\prec 1,$ (SA.6.27)
@@
By (SA.6.25), $\lvert\sigma_{n}-\tilde{\sigma}_{n}\rvert\prec 1$, where
$\tilde{\sigma}_{n}^{2}=\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}\text{Cov}({\bf E}[Z_{i}\mid\mathcal{F}_{i}(r_{n})],{\bf E}[Z_{j}\mid\mathcal{F}_{j}(r_{n})])A_{ij}(1)$ (SA.6.28)
@@
$\hat{\sigma}^{2}(1)=\tilde{\sigma}_{n}^{2}+\mathcal{B}_{n}+o_{p}(1).$ (SA.6.29)
@@
$\lvert\hat{\sigma}^{2}(1)-\check{\sigma}^{2}(1)\rvert\leq\frac{2k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\lvert Z_{i}\rvert\sum_{j\in\mathcal{N}_{n}}A_{ij}(1)\max_{k}\lvert\hat{Z}_{k}-Z_{k}\rvert\\ +\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}A_{ij}(1)\max_{k}(\hat{Z}_{k}-Z_{k})^{2}.$ (SA.6.30)
@@
which is $\prec 1$ by the proof of Theorem 2. Hence, (SA.6.30)1precedesitalic-(SA.6.30italic-)$\prec 1$.

### v4

@@ 8 Conclusion
where $\mathcal{N}(i,-1)\equiv\varnothing$. As $r$ increases, $J(r,C_{j})$ moves away from the boundary and towards the interior. If $i\in J(r,C_{j})$, then $\mathcal{N}(i,r)\subseteq C_{j}$, so for such $i$, by Assumptions 3 and 4,
@@
$\leq 2c\,\min\{r^{-\gamma},1\}$
for any $t\in\{0,1\}$. Then
@@
$\precsim\frac{1}{n}\sum_{j=1}^{k_{n}}\left(\lvert\mathcal{B}(C_{j})\rvert+\sum_{r=0}^{L_{n}}\sum_{i\in J(r,C_{j})}\min\{r^{-\gamma},1\}\right)$
$\precsim\underbrace{\frac{k_{n}}{n}\xi_{n}\left(\frac{n}{k_{n}\xi_{n}}\right)^{\frac{d-1}{d}}}_{\text{boundary bias}}+\frac{k_{n}}{n}\sum_{r=0}^{L_{n}}\xi_{n}\max\{(L_{n}-r-1)^{d-1},1\}\min\{r^{-\gamma},1\}$
$\precsim(k_{n}\xi_{n}/n)^{1/d}+\frac{k_{n}\xi_{n}}{n}L_{n}^{d-1}\sum_{r=0}^{\infty}\min\{r^{-\gamma},1\}\precsim(k_{n}\xi_{n}/n)^{1/d}.$
@@
Let $d\geq 1$, and fix any sequence $\{\xi_{n}\}_{n\in\mathbb{N}}$ such that $1\precsim\xi_{n}\prec n$. Let $\rho$ be the sup norm and $\mathcal{N}_{n}=\{\xi_{n}^{-1/d}x\colon x\in\mathbb{Z}^{d}\}\cap B(\bm{0},\mathcal{R}_{n})$ where $B(\bm{0},\mathcal{R}_{n})$ is a (hyper)cube with radius $\mathcal{R}_{n}\in\mathbb{Z}$ and $\mathcal{R}_{n}\sim(n/\xi_{n})^{1/d}$. When $\xi_{n}\sim 1$, the observed units are positioned within a cube of radius $\sim n^{1/d}$ containing $\sim n$ units positioned on the integer lattice. When $\xi_{n}\succ 1$, we shrink this region towards the origin by a factor $\xi_{n}^{-1/d}$. This results in the same asymptotic order of number of units $\lvert\mathcal{N}_{n}\rvert\sim n$, but the volume of the cube is reduced to $\sim\mathcal{R}_{n}^{d}\sim n/\xi_{n}$, resulting in a density of $\xi_{n}$.
@@
$\lvert\mathcal{N}(i,r+\xi_{n}^{-1/d})\backslash\mathcal{N}(i,r)\rvert\in[\min\{C_{1}^{-1}\xi_{n}^{(d-1)/d}r^{d-1},n\},\max\{C_{1}\xi_{n}^{(d-1)/d}r^{d-1},1\}].$
Letting $[c]$ denote rounding $c$ to the nearest integer, this implies
@@
$\leq 1+C_{1}d^{-1}\xi_{n}((r+\xi_{n}^{-1/d}[\xi_{n}^{1/d}])^{d}-r^{d}),$ (SA.6.17)
which is bounded by a constant times $\xi_{n}r^{d-1}$. This verifies Assumption 2.
Verifying Assumption 5. Let $1\prec k_{n}\prec n/\xi_{n}$. For the remainder of the proof, suppose the clusters are generated by $k_{n}$-medoids which by Theorem 1 satisfy Assumption 5. Also let $n$ be sufficiently large that $L_{n}$ in the assumption exceeds two.
@@
$Y_{i}(\bm{d})=d_{i}\frac{\sum_{j\in\mathcal{N}_{n}}d_{j}\bm{1}\{\rho(i,j)\leq 2\}}{\sum_{j\in\mathcal{N}_{n}}\bm{1}\{\rho(i,j)\leq 2\}},$
which is unit $i$’s treatment times the fraction of treated units in $i$’s 2-neighborhood. The sum in the denominator is always at least one since it includes $i$ itself, so Assumption 1 holds. Assumption 3 holds because potential outcomes only depend on 2-neighborhood treatments.
Reduction to HT. Notice $Y_{i}(\bm{0})=0$. Since $Q\in\{D,T,O\}$ and $p_{0}=0$, $T_{0i}^{+}Y_{i}=0$. Then
@@
Bias. Let $\beta=(p_{1}-p_{1}q)\bm{1}\{d_{1}=1\}+p_{1}(p_{1}-p_{1}q)\bm{1}\{d_{1}=\varnothing\}$, which is strictly positive by assumption. We have
$\lvert{\bf E}[\tilde{\theta}^{+}]-\theta^{*}\rvert=\bigg{|}\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}\big{(}{\bf E}[Y_{i}\mid T_{1i}^{+}=1]-{\bf E}^{*}_{p_{1}}[Y_{i}(d_{1},\bm{D}_{-i})]\big{)}\bigg{|}\\ =\beta\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}\frac{\sum_{j\in\mathcal{N}_{n}\backslash C_{c(i)}}\bm{1}\{\rho(i,j)\leq 2\}}{\sum_{j\in\mathcal{N}_{n}}\bm{1}\{\rho(i,j)\leq 2\}}.$ (SA.6.18)
@@
~$\tilde{\mathcal{B}}(C)=\big{\{}i\in C\colon\lvert\mathcal{N}(i,1)\cap(\{\xi_{n}^{-1/d}x\colon x\in\mathbb{Z}^{d}\}\backslash C)\rvert\geq 1\big{\}},$
the set of units whose $1$-neighborhoods intersect a point outside the set. Then
(SA.6.18)$\geq\beta\frac{1}{n}\sum_{j=1}^{k_{n}}\sum_{i\in\tilde{\mathcal{B}}(C_{j})\backslash\tilde{\mathcal{B}}(\mathcal{N}_{n})}\frac{\sum_{\ell\in\mathcal{N}_{n}\backslash C_{j}}\bm{1}\{\rho(i,\ell)\leq 2\}}{\sum_{\ell\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell)\leq 2\}}.$ (SA.6.19)
By construction of $\mathcal{N}_{n}$,
$\min_{j=1,\ldots,k_{n}}\min_{i\in\tilde{\mathcal{B}}(C_{j})\backslash\tilde{\mathcal{B}}(\mathcal{N}_{n})}\frac{\sum_{\ell\in\mathcal{N}_{n}\backslash C_{j}}\bm{1}\{\rho(i,\ell)\leq 2\}}{\sum_{\ell\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell)\leq 2\}}\succsim 1.$
@@
Let $m_{j}$ denote the medoid of a cluster $C_{j}$. By Assumption 5(a), units in ~$\tilde{\mathcal{B}}(C_{j})$ are at least distance $\lfloor L_{n}\rfloor$ from $m_{j}$. Define the “contour set”
$\tilde{J}(r,C_{j})=\mathcal{N}(m_{j},r)\backslash\mathcal{N}(m_{j},r-1).$
@@
$\sum_{\ell\in\mathcal{N}_{n}}\sum_{m\in\mathcal{N}_{n}}\frac{\bm{1}\{\rho(i,\ell)\leq 2\}}{\sum_{\ell^{\prime}\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell^{\prime})\leq 2\}}\frac{\bm{1}\{\rho(j,m)\leq 2\}}{\sum_{m^{\prime}\in\mathcal{N}_{n}}\bm{1}\{\rho(j,m^{\prime})\leq 2\}}\\ \times(p_{1}^{+})^{-2}\text{Cov}\big{(}T_{1i}^{+}D_{i}D_{\ell},T_{1j}^{+}D_{j}D_{m}\big{)}.$
@@
$\text{Cov}(W_{c(i)}D_{i}D_{\ell},W_{c(j)}D_{j}D_{m})\\ =\underbrace{{\bf E}[W_{c(i)}\text{Cov}(D_{i}D_{\ell},D_{j}D_{m}\mid W_{c(i)})]}_{\geq 0}+\text{Cov}({\bf E}[D_{i}D_{\ell}\mid W_{c(i)}],{\bf E}[D_{j}D_{m}\mid W_{c(i)}]).$
@@
$\text{Var}\left(\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}Z_{i}^{+}\right)\geq\alpha\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in C_{c(i)}}\sum_{\ell\in C_{c(i)}\backslash\{i\}}\frac{\bm{1}\{\rho(i,\ell)\leq 2\}}{\sum_{\ell^{\prime}\in\mathcal{N}_{n}}\bm{1}\{\rho(i,\ell^{\prime})\leq 2\}}\\ \times\sum_{m\in C_{c(i)}\backslash\{j\}}\frac{\bm{1}\{\rho(j,m)\leq 2\}}{\sum_{m^{\prime}\in\mathcal{N}_{n}}\bm{1}\{\rho(j,m^{\prime})\leq 2\}}\succsim\frac{1}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\lvert C_{c(i)}\rvert.$
@@
By Assumption 3,
$\lvert Y_{i}(\bm{D})-{\bf E}[Y_{i}(\bm{D})\mid\mathcal{F}_{i}(r)]\rvert=\lvert Y_{i}(\bm{D})\pm Y_{i}(\bm{D}_{\mathcal{N}(i,r)},\bm{0}_{-\mathcal{N}(i,r)})\\ -{\bf E}[Y_{i}(\bm{D})\pm Y_{i}(\bm{D}_{\mathcal{N}(i,r)},\bm{0}_{-\mathcal{N}(i,r)})\mid\mathcal{F}_{i}(r)]\rvert\leq 2c\min\{r^{-\gamma},1\}$
since ${\bf E}[Y_{i}(\bm{D}_{\mathcal{N}(i,r)},\bm{0}_{-\mathcal{N}(i,r)})\mid\mathcal{F}_{i}(r)]=Y_{i}(\bm{D}_{\mathcal{N}(i,r)},\bm{0}_{-\mathcal{N}(i,r)})$. Then by Lemma SA.5.3, there exists $c_{n}\precsim 1$ such that for all $i$,
$\lvert Z_{i}-{\bf E}[Z_{i}\mid\mathcal{F}_{i}(r)]\rvert\leq c_{n}\min\{r^{-\gamma},1\}\quad\text{for all}\quad r\geq r_{n}.$ (SA.6.23)
By Lemma SA.5.4, $\max_{i}\lvert\Lambda_{i}\rvert\precsim n/k_{n}$, so (SA.6.11) and (SA.6.23) imply
$\lvert[P1]\rvert\precsim(c_{n}r_{n}^{-\gamma})^{2}\precsim(k_{n}\xi_{n}/n)^{2\gamma/d},$
@@
Turning to $\lvert[P2]\rvert$, fix $i,j$ such that $j\not\in\Lambda_{i}$. By (SA.6.23), there exists $c_{n}\precsim 1$ such that for all such $i,j$,
$\lvert{\bf E}[B_{i}B_{j}]\rvert\leq c_{n}r_{n}^{-\gamma}.$
@@
$B_{i}-\underbrace{{\bf E}[B_{i}\mid\mathcal{F}_{i}(\rho(i,j)/2-\bar{R})]}_{X_{i}}=Z_{i}-{\bf E}[Z_{i}\mid\mathcal{F}_{i}(\rho(i,j)/2-\bar{R})]$
since $\mathcal{N}(i,\rho(i,j)/2-\bar{R})\supseteq\mathcal{N}(i,r_{n})$. Notice ${\bf E}[X_{i}]=0$; $\max_{i}\lvert X_{i}\rvert\precsim 1$ by Assumption 1 and Lemma SA.5.3; and $X_{i}\perp\!\!\!\perp X_{j}$ by Assumption 4 since $\mathcal{N}(i,\rho(i,j)/2-\bar{R})$ and $\mathcal{N}(j,\rho(i,j)/2-\bar{R}))>2\bar{R}$ are separated by a distance of at least $2\bar{R}$, which is an upper bound on the “diameter” of any cluster. Applying (SA.6.23) with $\rho(i,j)/2-\bar{R}$ in place of $r_{n}$, there exists $c_{n}^{\prime}\precsim 1$ such that for any $i,j$,
$\lvert{\bf E}[B_{i}B_{j}]\rvert\leq\lvert{\bf E}[X_{i}X_{j}]\rvert+\lvert{\bf E}[X_{i}(B_{j}-X_{j})]\rvert\\ +\lvert{\bf E}[(B_{i}-X_{i})X_{j}]\rvert+\lvert{\bf E}[(B_{i}-X_{i})(B_{j}-X_{j})]\rvert\leq c_{n}^{\prime}(\rho(i,j)/2-\bar{R})^{-\gamma}.$
@@
$\lvert[P2]\rvert\leq\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\not\in\Lambda_{i}}\lvert{\bf E}[B_{i}B_{j}]\rvert\\ \leq k_{n}\frac{c_{n}^{\prime}}{n^{2}}\sum_{\ell=1}^{k_{n}}\sum_{i\in C_{\ell}}\sum_{s=\lfloor 2r_{n}\rfloor}^{\infty}\sum_{j\not\in\Lambda_{i}}\bm{1}\{\rho(i,j)\in[s,s+1)\}\\ \times\big{(}r_{n}^{-\gamma}\bm{1}\{s\leq 4\bar{R}\}+(s/2-\bar{R})^{-\gamma}\bm{1}\{s>4\bar{R}\}\big{)}.$
@@
$\lvert\sigma_{n}-\tilde{\sigma}_{n}\rvert\leq\text{Var}\left(\sqrt{k_{n}}\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}(Z_{i}-{\bf E}[Z_{i}\mid\mathcal{F}_{i}(r_{n})])\right)^{1/2}\prec 1$ (SA.6.24)
@@
$\tilde{\sigma}_{n}^{-1}\sqrt{k_{n}}\frac{1}{n}\sum_{i\in\mathcal{N}_{n}}\big{(}{\bf E}[Z_{i}\mid\mathcal{F}_{i}(r_{n})]-{\bf E}[Z_{i}]\big{)}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathcal{N}(0,1)$ (SA.6.25)
@@
so (SA.6.25) follows from his (3.8). This completes the proof of (5)
@@
$\sqrt{k_{n}}(\bar{\theta}-\theta^{*})\precsim\sqrt{k_{n}}(k_{n}\xi_{n}/n)^{\gamma/d}\prec\sqrt{k_{n}}(n/\xi_{n})^{-\frac{\gamma}{2\gamma+d}}\prec 1,$ (SA.6.26)
@@
By (SA.6.24), $\lvert\sigma_{n}-\tilde{\sigma}_{n}\rvert\prec 1$, where
$\tilde{\sigma}_{n}^{2}=\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}\text{Cov}({\bf E}[Z_{i}\mid\mathcal{F}_{i}(r_{n})],{\bf E}[Z_{j}\mid\mathcal{F}_{j}(r_{n})])A_{ij}(1)$ (SA.6.27)
@@
$\hat{\sigma}^{2}(1)=\tilde{\sigma}_{n}^{2}+\mathcal{B}_{n}+o_{p}(1).$ (SA.6.28)
@@
$\lvert\hat{\sigma}^{2}(1)-\check{\sigma}^{2}(1)\rvert\leq\frac{2k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\lvert Z_{i}\rvert\sum_{j\in\mathcal{N}_{n}}A_{ij}(1)\max_{k}\lvert\hat{Z}_{k}-Z_{k}\rvert\\ +\frac{k_{n}}{n^{2}}\sum_{i\in\mathcal{N}_{n}}\sum_{j\in\mathcal{N}_{n}}A_{ij}(1)\max_{k}(\hat{Z}_{k}-Z_{k})^{2}.$ (SA.6.29)
@@
which is $\prec 1$ by the proof of Theorem 2. Hence, (SA.6.29)1precedesitalic-(SA.6.29italic-)$\prec 1$.